. Order: Acyclovir 24 g over 24 hour IV

            Supply: Acyclovir 500 mg in 100 mL D5W

What is the rate?_____________

. Order: Acyclovir 24 g over 24 hour IV

            Supply: Acyclovir 500 mg in 100 mL D5W

What is the rate?_____________

. Order: Acyclovir 24 g over 24 hour IV

            Supply: Acyclovir 500 mg in 100 mL D5W

What is the rate?_____________

. Order: Acyclovir 24 g over 24 hour IV

            Supply: Acyclovir 500 mg in 100 mL D5W

What is the rate?_____________
Answer #1

ANSWER.

* Calculation of rate.

          Rate (ml/hr) = Volume(ml) / Time(hr).

    Calculation of volume(ml)

          Volume(ml) = Desired dose/Dose in hand * Quantity.

     Here

         Desired dose(D) = 24000 mg (24g)

         Dose in hand (H) = 500 mg.

               Quantity (Q) = 100 ml.

As per above formula

      Volume(ml) of drug required = D/H * Q

                                                    = 24000 / 500 * 100

                                                    = 48 * 100

                                                    = 4800 ml.

                          Time (hr) = 24 hr.

           Flow rate = Volume(ml) /Time(hr)

                           = 4800 / 24

                           = 200 ml/hr.